Let X and Y be normed spaces, T∈B(X,Y) and (xn) a sequence in X. If xnx0, show that TxnTx0.

1
Expert's answer
2021-09-28T01:17:42-0400

ANSWER

Since "B(X,Y)" is a space of linear bounded operators in which norma can be given "\left\| T \right\| =sup\left\{ { \left\| Tx \right\| }_{ Y } \right| x\in X,\ { \left\| x \right\| }_{ X }=1\} \quad ," then for "x" non equal to zero "{ \left\| T\left( \frac { x }{ \left\| x \right\| } \right) \right\| }_{ Y }\le \ { \left\| T \right\| }_{ B(\{ ,Y) }" . The mapping "T" is a linear operator , so "T\left( \frac { x }{ \left\| x \right\| } \right) =\frac { 1 }{ \left\| x \right\| } Tx" . Therefore, for all "x\in X" the inequality

. . "{ \left\| Tx \right\| }_{ Y }\le \left\| T \right\| \cdot { \left\| x \right\| }_{ X }" (1)

is true.

"T\left( { x }_{ n }-{x}_0\ \right) =T{ x }_{ n }-T{x}_0" , "{ \left\| T{ x }_{ n }-T{x}_0 \right\| }_{ Y }={ \left\| T\left( { x }_{ n }-{x}_0\ \right) \right\| }_{ Y }" . From (1) it follows

"{ \left\| T\left( { x }_{ n }-{ x }_{ 0 }\ \right) \right\| }_{ Y }\le \left\| T \right\| \cdot { \left\| { x }_{ n }-{ { x }_{ 0 } } \right\| }_{ X }" or

"0\le { \left\| T{ x }_{ n }-T{ x }_{ 0 } \right\| }_{ Y }\le \left\| T \right\| \cdot { \left\| { x }_{ n }-{ { x }_{ 0 } } \right\| }_{ X }" (2)

By definition: "{ x }_{ n }\rightarrow { { x }_{ 0 } }\ \Leftrightarrow \ { \left\| { x }_{ n }-{ { x }_{ 0 } } \right\| }_{ X }\rightarrow 0" .

By the Sandwich Theorem from (2) it follows "{ \left\| T{ x }_{ n }-T{ x }_{ 0 } \right\| }_{ Y }\rightarrow 0\quad"or "T{ x }_{ n }\rightarrow T{ { x }_{ 0 } }" .